If the same number is added to both the numerators and denominator of the fraction 3/5, the value of the resulting fraction is 8/9. Find the number.
Answer by amoresroy(361)

Let x = the number
(3+x)/(5+x) = 8/9
Cross multiply to get
9((3+x) = 8(5+x)
27 + 9x = 40 + 8x
9x-8x = 40 - 27
x = 13